A $30,000 bond pays interest 6% annually on October 1st and matures on October 1, 2028. The bond is purchased on May 1, 2017 to yield 8% interest compounded annually. Compute the purchase price.
Answer #1

Rate = 8%
PMT = 30,000 * 6% = 1800
FV = 30000
Nper = 11+5/12

Purchase price can be calculated by using the following excel formula:
=PV(rate,nper,pmt,fv)
=PV(8%,11+5/12,-1800,-30000)
= $25,615.11

Purchase price = $25,615.11 or $25,615 (approx)

FEEL FREE TO ASK DOUBTS.